OpenStudy (anonymous):

What is the common ratio of this geometric sequence? 1 27 2 18 3 12 4 8 5 16/3 A. -9 B. 2/3 C. 3/2 D. 9

OpenStudy (anonymous):

can u see the pattern 18/27 =12/18 = 2/3

OpenStudy (anonymous):

i dont see it.. @ritez

OpenStudy (anonymous):

wait is it -9? @ritez

OpenStudy (anonymous):

no the ratio is 2/3 as 18/27 =12/18 =8/12 =2/3

OpenStudy (anonymous):

what about 1 50 2 30 3 18 4 10.8 5 6.48 A. -20 B. 0.6 C. 1.67 D. 20 @ritez

OpenStudy (anonymous):

30/50 = 18/30 can u figure out now?

OpenStudy (anonymous):

yess, thank you :) @ri

OpenStudy (anonymous):

@ritez

OpenStudy (anonymous):

yw:)